# Spec-Gloss Surfels and Normal–Diffuse Priors for Relightable Glossy Objects

### v1

@@ 3.3 Coarse-to-fine HDR Lighting
We represent scene illumination using a learnable HDR environment map $E$, parameterized as a cube mipmap. During training, the cube map provides efficient prefiltered lookups for the split-sum approximation based on [17].
Progressive Upsampling (PU). To mitigate the computational bottleneck of mipmap generation in the split-sum formulation, we adopt a coarse-to-fine training strategy. The cubemap is initialized at a low per-face resolution, $\text{Res}^{\text{init}}$, and progressively doubled at fixed iteration intervals until $\text{Res}_{\text{final}}$. At each step, the environment map is bilinearly interpolated, and the cube mipmaps are rebuilt down to $\text{Res}_{\text{min}}$, reducing early training cost, preventing lighting from absorbing material errors, and ensuring robust coarse-scale convergence before high-frequency lighting is introduced. At the same time, we avoid redundant smoothing operations as in prior works [36, 35, 29]. Formally, given $E^{k}\in\mathbb{R}^{6\times\text{Res}^{k}\times\text{Res}^{k}\times 3}$, an HDR cubemap at stage k=1,,Kk=1,...,K, we perform progressive upsampling of $E$, defined as $E^{k+1}=\mathcal{U}_{\text{r}_{k}\rightarrow\text{r}_{k+1}}(E^{k})$, where $\text{Res}_{\text{k+1}}=2\times\text{Res}^{\text{k}}$ and $\mathcal{U}$ denotes the billinear upsampling operation. Then we rebuild the prefiltered cube mip pyramid by successively downsampling from the highest resolution $\text{Res}^{\text{k}}$ to $\text{Res}^{\text{min}}$.
$\{M_{j}^{k}\}_{j}=\text{PMREM}(E^{k}),$ (6)
where PMREM stands for ”Prefiltered, Mipmapped Radiance Environment Map” and is an operation that successively downsamples the cubemap with average pooling until reaching the minimum resolution $\text{Res}^{\text{min}}$, thus generating a cube mipmap, as shown in Fig. 4.
@@
Negative-only Clipping (NOC). In reality, environment lights contribute to dynamic contrast, so its range should not be restricted. We enforce non-negativity on the HDR envmap and leave positives unbounded: $E\leftarrow\max(0,E_{\text{raw}})$ each iteration. Prior work clamps $E_{\text{raw}}\in[0,1]$ [12, 33], which underexposes lighting and deteriorates material-light disentanglement. NOC preserves HDR peaks required for sharp, bright specular reflections and avoids systematic underexposure, as shown in Fig. 5.
Figure 5: Negative-only clipping (NOC) preserves the specular highlights and improves illumination recovery.
@@ 3.4 Ambiguity Suppression
Figure 6: The normal prior (shown in camera coordinates) reduces the ambiguity caused by interreflections and improves reconstruction quality as seen in the predicted normals (shown in world coordinates).
We leverage StableNormal [34] to provide per-pixel estimates of surface normals. These normals act as soft constraints on the reconstructed G-buffer, encouraging consistent geometry (see Fig. 6) and sharper specular highlights. Specifically, the normal prior $\tilde{N}$ is used as pseudo-GT for the predicted normals $\hat{N}$ and the normals derived from the gradient of the surface depth $\hat{N}_{D}$. The normal prior regularization term is defined as
@@
Diffuse Color Prior. We incorporate a diffuse color prior based on StableDelight [25], a diffusion model that removes specular reflections from natural images, thus generating a powerful prior for the diffuse color of a scene, as demonstrated in Fig. 7. This prior provides a strong initialization, guiding the network toward realistic color decomposition in the early phase of training. While the prior offers useful guidance, its accuracy may be surpassed by the results obtained after scene optimization. To exploit its benefits while mitigating these drawbacks, we apply it only as a soft prior at the beginning of training and then disable it, allowing the model to optimize solely based on the RGB ground truth. The diffuse color prior term is defined as
$\mathcal{L}_{\text{dp}}=\lambda_{\text{dp}}(t)\,\|I_{\text{d}}-I_{\text{dp}}\|_{1},\quad\lambda_{d}(t)=\begin{cases}\lambda_{\text{dp}},&t\leq T_{\text{dp}},\\ 0,&t>T_{\text{dp}},\end{cases}$ (8)
@@
Figure 7: The diffuse prior steers the optimization of albedo and illumination into the right direction.
@@ 3.5 Composite Objective
Our objective loss function follows 2DGS [8] with an RGB reconstruction loss and a D-SSIM loss in sRGB space,
@@
where $\lambda$ is the balancing weight and is set to 0.2 by default. We also use the cosine distance loss $\mathcal{L}_{\text{n}}$ that enforces consistency between the predicted normals $\hat{N}$ and the normals derived from the depth gradients $\hat{N}_{D}$. Finally, we add the normal prior loss $\mathcal{L}_{\text{np}}$, the diffuse prior loss $\mathcal{L}_{\text{dp}}$, and the white-light regularization loss $\mathcal{L}_{\text{light}}$ used in [21, 27, 6]. The total objective loss function is defined as
@@ 4.1 Implementation Details
We train synthetic scenes for 50k iterations and real scenes for 20k iterations on an NVIDIA GeForce RTX 4090. Deferred shading is used throughout training. The indirect lighting stage begins at 20k iterations for synthetic scenes and 10k for real scenes. After this point, we extract an object mesh every 3k iterations to enable visibility estimation via ray tracing. Learning rates are set to 0.005 for $F_{0}$ and roughness, 0.0075 for albedo, and 0.01 for the environment lighting cubemap, each decayed by a factor of 0.1. All other trainable attributes follow the settings of 2DGS [8]. Loss weights are set to $\lambda=0.2$, $\lambda_{\text{n}}=0.05$, $\lambda_{\text{np}}=0.01$, $\lambda_{\text{dp}}=0.05$, and $\lambda_{\text{light}}=0.001$. The cutoff time for the diffuse prior is set as $T_{\text{dp}}=15k$, and the prior is disabled when it deteriorates performance in failure cases of StableDelight [25] e.g. mirror-like objects like teapot from the Glossy Synthetic dataset [21]. For PU, we start with a face resolution $\text{Res}^{\text{init}}=64$ and progressively upsample the cubemap faces every 1515k iterations until reaching $\text{Res}^{\text{final}}=512$. The cubemap is prefiltered in every iteration until the base resolution of $\text{Res}^{\text{min}}=16$ for approximating different levels of roughness from $r^{\text{min}}=0.02$ to $r^{\text{max}}=0.5$.
@@ 4.2 Comparison
Relighting. Tab. 1 and Fig. 9 present our relighting results. Relighting is performed using an HDR environment map unseen during training, converted from latlong format into a cubemap, then prefiltered across mip levels to approximate varying roughness levels. Our method directly uses the raw HDR radiance values, while other methods apply tone mapping as specified in their pipelines. We do not apply any rescaling based on GT relighted images, or any other post-processing step and report the raw relighted images generated, solely based on recovered material properties and the novel HDR environment map. Both quantitative and qualitative evaluations show that our approach achieves superior relighting performance on complex glossy objects. We also compare results across the three material parameterizations: SG, MR, and MRS, discussed in Sec. 3.2. Our method with the SG parameterization consistently outperforms the baselines and variants and achieves improved relighting fidelity thanks to more plausible material estimates. Fig. 8 shows the recovered material properties and diffuse/specular outputs for our method vs MSR-based Ref-Gaussian [33]. The latter tends to overestimate metallic as can be seen on the visor and stripes of the helmet and the surface of the coffee.
@@
Table 1: Relighting evaluation averaged over three previously unseen environment maps (corridor, golf, neon) on the Glossy Synthetic [21] scenes. Results for R3DG [6] and GS-2DGS [27] are based on [27]. The rest are obtained through the released codebases. We highlight the first, second, and third best results.
@@
GShader [9] 20.34 15.92 15.83 14.04 18.47 19.25 17.31
R3DG [6] 18.71 20.22 20.30 19.81 16.50 17.22 18.79
GS-2DGS [27] 18.95 21.78 18.82 17.88 17.55 18.96 18.99
3DGS-DR [12] 20.59 20.36 21.34 20.01 19.08 19.94 20.22
RefGauss [33] 19.87 21.58 20.45 20.20 21.35 21.45 20.82
@@
GShader [9] 0.900 0.868 0.828 0.783 0.875 0.883 0.856
R3DG [6] 0.840 0.839 0.862 0.903 0.819 0.798 0.818
GS-2DGS [27] 0.860 0.851 0.879 0.919 0.824 0.824 0.866
3DGS-DR [12] 0.886 0.894 0.875 0.862 0.883 0.906 0.884
RefGaass [33] 0.888 0.919 0.877 0.870 0.912 0.894 0.893
@@
GShader [9] 0.114 0.104 0.098 0.156 0.138 0.115 0.121
R3DG [6] - - - - - - -
GS-2DGS [27] - - - - - - -
3DGS-DR [12] 0.103 0.097 0.083 0.115 0.110 0.080 0.098
RefGauss [33] 0.104 0.070 0.079 0.110 0.087 0.087 0.090
@@
Figure 9: Relighting comparison on three scenes in the Glossy Synthetic [21] dataset, with the used environment map on the left.
Reconstruction and Lighting Recovery. Tab. 2 reports geometry reconstruction quality on the Shiny Synthetic dataset [30], environment map quality across two synthetic datasets [30, 21], average training time, and rendering speed. Geometry is evaluated via MAE of surface normals, and illumination via LPIPS on extracted HDR latlong maps. Our method consistently outperforms Gaussian Splatting baselines in both tasks, with minimal added training cost, showcasing the benefit of our contributions. Qualitative examples of recovered environment maps are shown in Fig. 10, with more results in the supplementary material.
@@
ENVIDR [19] 2.74 0.501 5.84 1
GShader [9] 7.00 0.635 0.48 28
3DGS-DR [12] 2.62 0.584 0.35 251
Ref-GS [37] 2.21 - 0.64 45
RefGauss [33] 2.15 0.560 0.58 122
@@
Figure 11: Novel View Synthesis on three scenes from [30, 21]. Our method reconstructs finer details and preserves sharper reflections and specular highlights.
@@
Method Rel. Shiny Synthetic [30] Glossy Synthetic [21] Shiny Real [30]
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Ref-NeRF [30] 33.13 0.961 0.080 27.50 0.927 0.100 23.62 0.646 0.239
3DGS [11] 30.36 0.947 0.084 26.50 0.916 0.092 23.85 0.660 0.230
GShader[9] 31.97 0.958 0.067 27.54 0.922 0.087 23.46 0.647 0.257
ENVIDR [19] 33.46 0.979 0.046 29.58 0.952 0.057 23.00 0.606 0.332
3DGS-DR[12] 34.09 0.971 0.057 30.22 0.953 0.061 23.99 0.664 0.229
Ref-GS[37] 34.80 0.973 0.056 30.58 0.957 0.058 24.44 0.657 0.225
RefGauss 1 [33] 35.04 0.970 0.056 30.92 0.964 0.047 23.98 0.663 0.278
@@
    For RefGauss, although higher scores were reported on Shiny Real we were unable to reproduce them. We therefore report results obtained using the authors’ released code.
Table 4: Ablation study on the Glossy Synthetic [30] dataset comparing relighting, novel view synthesis quality, and training time. PU stands for Progressive Upsampling, Priors denotes the normal and diffuse priors, and NOC stands for Negative-only Clipping.
@@ 4.3 Ablation Study
We ablate progressive upsampling (PU), StableNormal [34] and StableDelight [25] priors (Priors), and negative only clipping (NOC) on Glossy Synthetic. As shown in Tab. 4, PU cuts training time from 112 to 61 minutes with no NVS drop and a small relighting gain. Priors improve NVS by disambiguating geometry and material properties. Relighting changes little with Priors alone because lighting remains clipped, and specular cues are bounded. Combining Priors with NOC improves relighting dramatically and also boosts NVS. It can be concluded that Priors reduce ambiguity between geometry, materials and light, and NOC preserves HDR peaks needed for accurate specular IBL. A secondary benefit of PU is that it makes high-resolution envmaps feasible at reasonable training time. As the last ablation row shows, using a high envmap resolution yields a modest lift in relighting and NVS, so the gain is incremental. Our relighting improvements arise mainly from the Spec-Gloss parameterization, PU, Priors, and NOC rather than envmap resolution. Even at low resolution, we are on par with Ref-Gaussian [33] on novel view synthesis, but strongly outperform it on relighting at roughly equal training time.

### v2

@@ 3.3 Coarse-to-fine HDR Lighting
We represent scene illumination using a learnable HDR environment map $E$, parameterized as a cube mipmap. During training, the cube map provides efficient prefiltered lookups for the split-sum approximation based on [Laine2020diffrast].
Progressive Upsampling (PU). To mitigate the computational bottleneck of mipmap generation in the split-sum formulation, we adopt a coarse-to-fine training strategy. The cubemap is initialized at a low per-face resolution, $\text{Res}^{\text{init}}$, and progressively doubled at fixed iteration intervals until $\text{Res}_{\text{final}}$. At each step, the environment map is bilinearly interpolated, and the cube mipmaps are rebuilt down to $\text{Res}_{\text{min}}$, reducing early training cost, preventing lighting from absorbing material errors, and ensuring robust coarse-scale convergence before high-frequency lighting is introduced. At the same time, we avoid redundant smoothing operations from prior works [zhang2021nerfactor, zhang2021physg, Ummenhofer2024OWL]. Formally, given $E^{k}\in\mathbb{R}^{6\times\text{Res}^{k}\times\text{Res}^{k}\times 3}$, an HDR cubemap at stage k=1,,Kk=1,...,K, we perform progressive upsampling of $E$, defined as $E^{k+1}=\mathcal{U}_{\text{r}_{k}\rightarrow\text{r}_{k+1}}(E^{k})$, where $\text{Res}_{\text{k+1}}=2\times\text{Res}^{\text{k}}$ and $\mathcal{U}$ denotes the bilinear upsampling operation. Then we rebuild the prefiltered cube mip pyramid $\{M_{j}^{k}\}_{j}=\text{PMREM}(E^{k})$, where PMREM stands for ”Prefiltered, Mipmapped Radiance Environment Map” and is an operation that successively downsamples the cubemap with average pooling until reaching the minimum resolution $\text{Res}^{\text{min}}$ from the highest resolution $\text{Res}^{\text{k}}$, thus generating a cube mipmap, as shown in Fig. 4.
@@
Negative-only Clipping (NOC). In reality, environment lights contribute to dynamic contrast, so its range should not be restricted. We enforce non-negativity on the HDR envmap and leave positives unbounded: $E\leftarrow\max(0,E_{\text{raw}})$ each iteration. Prior work clamps $E_{\text{raw}}\in[0,1]$ [ye2024gsdr, yao2025refgaussian], which underexposes lighting and deteriorates material-light disentanglement. NOC preserves HDR peaks required for sharp, bright specular reflections and avoids systematic underexposure, as shown in Fig. 5.
Figure 5: Negative-only clipping (NOC) preserves specular highlights and improves illumination recovery. “w/o NOC” denotes standard clipping to the [0,1] range, where both negative values and values above 1 are clipped.
@@ 3.4 Ambiguity Suppression
Figure 6: The surface normal prior reduces ambiguity from view-dependent effects such as interreflections, yielding more accurate predicted normals and improved reconstruction quality.
We leverage StableNormal [ye2024stablenormal] to provide per-pixel estimates of surface normals. These normals act as soft constraints on the reconstructed G-buffer, encouraging consistent geometry (see Fig. 6) and sharper specular highlights. Specifically, the normal prior $\tilde{N}$ is used as pseudo-GT for the predicted normals $\hat{N}$ and the normals derived from the gradient of the surface depth $\hat{N}_{D}$. The normal prior regularization term is defined as
@@
Diffuse Color Prior. We incorporate a diffuse color prior based on StableDelight [stabledelight2025], a diffusion model that removes specular reflections from natural images, thus generating a powerful prior for the diffuse color of a scene, as demonstrated in Fig. 7. This prior provides a strong initialization, guiding the network toward realistic color decomposition in the early phase of training. While the prior offers useful guidance, its accuracy may be surpassed by the results obtained after scene optimization. To exploit its benefits while mitigating these drawbacks, we apply it only as a soft prior at the beginning of training and then disable it, allowing the model to optimize solely based on the RGB ground truth. The diffuse color prior term is defined as
$\mathcal{L}_{\text{dp}}=\lambda_{\text{dp}}(t)\,\|I_{\text{d}}-I_{\text{dp}}\|_{1},\quad\lambda_{dp}(t)=\begin{cases}\lambda_{\text{dp}},&t\leq T_{\text{dp}},\\ 0,&t>T_{\text{dp}},\end{cases}$ (6)
@@
Figure 7: The diffuse color prior reduces ambiguity between albedo and illumination and improves their disentanglement.
@@ 3.5 Composite Objective
Our objective loss function follows 2DGS [huang20242dgs] with an RGB reconstruction loss and a D-SSIM loss in sRGB space,
@@
where $\lambda$ is the balancing weight and is set to 0.2 by default. We also use the cosine distance loss $\mathcal{L}_{\text{n}}$ that enforces consistency between the predicted normals $\hat{N}$ and the normals derived from the depth gradients $\hat{N}_{D}$. Finally, we add the normal prior loss $\mathcal{L}_{\text{np}}$, the diffuse prior loss $\mathcal{L}_{\text{dp}}$, and the white-light regularization loss $\mathcal{L}_{\text{light}}$ used in [liu2023nero, tong2025gs2dgs, R3DG2024]. The total objective loss function is defined as
@@ 4.1 Implementation Details
We train synthetic scenes for 50k iterations and real scenes for 20k iterations on an NVIDIA GeForce RTX 4090. Deferred shading is used throughout training. The indirect lighting stage begins at 20k iterations for synthetic scenes and 10k for real scenes. After this point, we extract an object mesh every 3k iterations to enable visibility estimation via ray tracing. Learning rates are set to 0.005 for $F_{0}$ and roughness, 0.0075 for albedo, and 0.01 for the environment lighting cubemap, each decayed by a factor of 0.1. All other trainable attributes follow the settings of 2DGS [huang20242dgs]. Loss weights are set to $\lambda=0.2$, $\lambda_{\text{n}}=0.05$, $\lambda_{\text{np}}=0.01$, $\lambda_{\text{dp}}=0.05$, and $\lambda_{\text{light}}=0.001$. The cutoff time for the diffuse prior is set as $T_{\text{dp}}=15k$, and the prior is disabled when it deteriorates performance in failure cases of StableDelight [stabledelight2025] e.g. mirror-like objects such as teapot from the Glossy Synthetic dataset [liu2023nero]. For PU, we start with a face resolution $\text{Res}^{\text{init}}=64$ and progressively upsample the cubemap faces every 1515k iterations until reaching $\text{Res}^{\text{final}}=512$. The cubemap is prefiltered in every iteration until the base resolution of $\text{Res}^{\text{min}}=16$ for approximating different levels of roughness from $r^{\text{min}}=0.02$ to $r^{\text{max}}=0.5$.
@@ 4.2 Comparison
Relighting. Tab. 1 and Fig. 9 present our relighting results. Relighting is performed using an HDR environment map unseen during training, converted from latlong format into a cubemap, then prefiltered across mip levels to approximate varying roughness levels. Our method directly uses the raw HDR radiance values, while other methods apply tone mapping as specified in their pipelines. We do not apply any rescaling based on GT relighted images, or any other post-processing step and report the raw relighted images generated, solely based on recovered material properties and the novel HDR environment map. Both quantitative and qualitative evaluations show that our approach achieves superior relighting performance on complex glossy objects. We also compare results across the three material parameterizations: SG, MR, and MRS, discussed in Sec. 3.2. Our method with the SG parameterization consistently outperforms the baselines and variants and achieves improved relighting fidelity thanks to more plausible material estimates. Fig. 8 shows the recovered material properties and diffuse/specular outputs for our method vs MRS-based Ref-Gaussian [yao2025refgaussian]. The latter tends to overestimate metallic as can be seen on the visor and stripes of the helmet and the surface of the coffee.
@@
Table 1: Relighting evaluation averaged over three previously unseen environment maps (corridor, golf, neon) on the Glossy Synthetic [liu2023nero] scenes. Results for R3DG [R3DG2024] and GS-2DGS [tong2025gs2dgs] are based on [tong2025gs2dgs]. The rest are obtained through the released codebases. We highlight the first, second, and third best results.
@@
GShader [jiang2024gshader] 20.34 15.92 15.83 14.04 18.47 19.25 17.31
R3DG [R3DG2024] 18.71 20.22 20.30 19.81 16.50 17.22 18.79
GS-2DGS [tong2025gs2dgs] 18.95 21.78 18.82 17.88 17.55 18.96 18.99
3DGS-DR [ye2024gsdr] 20.59 20.36 21.34 20.01 19.08 19.94 20.22
RefGauss [yao2025refgaussian] 19.87 21.58 20.45 20.20 21.35 21.45 20.82
@@
GShader [jiang2024gshader] 0.900 0.868 0.828 0.783 0.875 0.883 0.856
R3DG [R3DG2024] 0.840 0.839 0.862 0.903 0.819 0.798 0.818
GS-2DGS [tong2025gs2dgs] 0.860 0.851 0.879 0.919 0.824 0.824 0.866
3DGS-DR [ye2024gsdr] 0.886 0.894 0.875 0.862 0.883 0.906 0.884
RefGauss [yao2025refgaussian] 0.888 0.919 0.877 0.870 0.912 0.894 0.893
@@
GShader [jiang2024gshader] 0.114 0.104 0.098 0.156 0.138 0.115 0.121
R3DG [R3DG2024] - - - - - - -
GS-2DGS [tong2025gs2dgs] - - - - - - -
3DGS-DR [ye2024gsdr] 0.103 0.097 0.083 0.115 0.110 0.080 0.098
RefGauss [yao2025refgaussian] 0.104 0.070 0.079 0.110 0.087 0.087 0.090
@@
Figure 9: Relighting comparison on three scenes in the Glossy Synthetic [liu2023nero] dataset, with the used environment map on the left.
Reconstruction and Lighting Recovery. Tab. 2 reports geometry reconstruction quality on the Shiny Synthetic dataset [verbin2022refnerf], environment map quality across two synthetic datasets [verbin2022refnerf, liu2023nero], average training time, and rendering speed. Geometry is evaluated via MAE of surface normals, and illumination via LPIPS on extracted HDR latlong maps. Our method consistently outperforms Gaussian Splatting baselines in both tasks, with minimal added training cost, showcasing the benefit of our contributions. Qualitative examples of recovered environment maps are shown in Fig. 10, with more results in the supplementary material.
@@
ENVIDR [liang2023envidr] 2.74 0.501 5.84 1
GShader [jiang2024gshader] 7.00 0.635 0.48 28
3DGS-DR [ye2024gsdr] 2.62 0.584 0.35 251
Ref-GS [zhang2024refgs] 2.21 - 0.64 45
RefGauss [yao2025refgaussian] 2.15 0.560 0.58 122
@@
Figure 11: Novel View Synthesis on three scenes from [verbin2022refnerf, liu2023nero]. Our method reconstructs finer details and preserves sharper reflections and specular highlights.
@@
Method Rel. Shiny Synthetic [verbin2022refnerf] Glossy Synthetic [liu2023nero] Shiny Real [verbin2022refnerf]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Ref-NeRF [verbin2022refnerf] 33.13 0.961 0.080 27.50 0.927 0.100 23.62 0.646 0.239
3DGS [kerbl20233dgs] 30.36 0.947 0.084 26.50 0.916 0.092 23.85 0.660 0.230
GShader [jiang2024gshader] 31.97 0.958 0.067 27.54 0.922 0.087 23.46 0.647 0.257
ENVIDR [liang2023envidr] 33.46 0.979 0.046 29.58 0.952 0.057 23.00 0.606 0.332
3DGS-DR [ye2024gsdr] 34.09 0.971 0.057 30.22 0.953 0.061 23.99 0.664 0.229
Ref-GS [zhang2024refgs] 34.80 0.973 0.056 30.58 0.957 0.058 24.44 0.657 0.225
RefGauss [yao2025refgaussian]1 35.04 0.970 0.056 30.92 0.964 0.047 23.98 0.663 0.278
@@
    RefGauss [yao2025refgaussian] reported higher scores on Shiny Real, but we were unable to reproduce them and instead report the obtained results.
Table 4: Ablation study on the Glossy Synthetic [verbin2022refnerf] dataset comparing relighting, novel view synthesis quality, and training time. PU stands for Progressive Upsampling, Priors denotes the normal and diffuse priors, and NOC stands for Negative-only Clipping.
@@ 4.3 Ablation Study
We ablate progressive upsampling (PU), diffuse and normal priors (Priors), and negative-only clipping (NOC) on scenes from the Glossy Synthetic dataset [liu2023nero]. As shown in Tab. 4, PU cuts training time from 112 to 61 minutes with no NVS drop and a small relighting gain, showing that a coarse-to-fine envmap schedule is beneficial. Priors improve NVS by disambiguating geometry and material properties, yielding cleaner normals and a more stable diffuse/specular split. Relighting changes little with Priors alone because lighting remains clipped and specular cues are bounded. Combining Priors with NOC dramatically improves relighting and also boosts NVS, as Priors reduce geometry-material-lighting ambiguity and NOC preserves the HDR peaks needed for accurate specular IBL. A secondary benefit of PU is that it makes high-resolution envmaps feasible at reasonable training time. As the last row shows, our final model already performs well at an envmap resolution of 128, and increasing the resolution to 512 yields only a modest lift in relighting and NVS, so the gain is incremental. Our relighting improvements primarily stem from the SG parameterization, PU, Priors, and NOC rather than envmap resolution. Even at low resolution, we are on par with Ref-Gaussian [yao2025refgaussian] on novel view synthesis, but strongly outperform it on relighting at comparable training time.